OpenStudy (anonymous):

find the rate, round to the nearest tenth of a percent. Use 360 days. p $3000 t 90 days Interest $105 I dont know where to begin

jimthompson5910 (jim_thompson5910):

Let I = interest P = principle r = rate t = time (in years) The formula for simple interest is then I = Prt We're given p $3000 t 90 days Interest $105 So I = 105, P = 3000 and t = 90/360 = 0.25 This means that I = Prt becomes 105 = 3000*r*0.25 I'll let you solve for r

OpenStudy (anonymous):

thanks i got 14%

jimthompson5910 (jim_thompson5910):

you got it, nice job